OpenStudy (anonymous):

Integrate this following seriously hard equation urgh! The equation is (Tanx^3 + tanx) / (2 secx^2 + tanx - 5) dx Please

OpenStudy (callisto):

Is it \((tanx)^3\) or \(tan(x^3)\)? Same question for secx^2..

OpenStudy (anonymous):

Its like the first one

OpenStudy (calculusfunctions):

OK! May I assist you?

hartnn (hartnn):

\(\large \int \frac{tan x (1+tan^2x)}{2(1+tan^2x)+tan x -5}dx=\int \frac{tan x (sec^2x)}{2(tan^2x)+tan x -3}dx \)

OpenStudy (calculusfunctions):

@hartnn exactly! but you're just giving her the solution rather then helping her. How is she supposed to do other similar problems then?

hartnn (hartnn):

i am sorry

hartnn (hartnn):

only one step, ok

OpenStudy (calculusfunctions):

Thank you! Now would you @hartnn like to teach her to arrive at the correct solution or may I? It would help if you were to delete the answers you already wrote, although it's probably too late now sine she already saw them. Man! teach. Don't show off!

OpenStudy (anonymous):

Yes yes im here! Please do!

hartnn (hartnn):

now i won't give solution

OpenStudy (calculusfunctions):

OK @fatinatikah did you see any part of @hartnn 's solution?

OpenStudy (calculusfunctions):

Thank you @hartnn I appreciate that!

OpenStudy (anonymous):

I only saw the first one. Are there more?

OpenStudy (calculusfunctions):

That's ok, but did you understand what @hartnn did because he is smart. He was right.

OpenStudy (anonymous):

Okay i do. So for the next step wat technique should i use then?

OpenStudy (calculusfunctions):

@fatinatikah are you there. If you are could you please reply to my question. Did you understand what he did?

OpenStudy (anonymous):

Yes yes i do. I want to kmow the next step. How do u integrate it then?

OpenStudy (calculusfunctions):

Not right away! First we need to simplify the fraction a bit. Could you show me what you think you saw as the first step. Then we'll take it from there.

OpenStudy (calculusfunctions):

If you understood it then you should be able to rewrite it, correct?

OpenStudy (anonymous):

The first step is using the identity tan x (tan^2 x + 1) and also for the other one right?

OpenStudy (calculusfunctions):

Yes that was the numerator. Now what is the denominator?

OpenStudy (anonymous):

So it will be tan x sec^2 x for the numerator right?

OpenStudy (calculusfunctions):

sec²x in the denominator reminds you of which Pythagorean trigonometric identity?

OpenStudy (anonymous):

Then for the denominator is 2 tan^2 x + tan x - 5

OpenStudy (calculusfunctions):

Yes the numerator is correct. Now can you tell me the correct denominator for that numerator in that step? Please.

OpenStudy (calculusfunctions):

There is something missing. Can you figure it out? Do you know what it is?

OpenStudy (calculusfunctions):

What does sec²x equal? Think Pythagorean trigonometric identity.

OpenStudy (anonymous):

Wait. Im sorry. I dont know. Im trying to figure it out

OpenStudy (calculusfunctions):

What are the three Pythagorean trigonometric identities? Do you know?

OpenStudy (calculusfunctions):

Sorry, take your time.

OpenStudy (anonymous):

Is it tan^2 x. + 1 = sec^2 x ?

OpenStudy (calculusfunctions):

Perfect! So if you replace the sec²x in the denominator with it's identity, how does the denominator simplify? Work it out with pencil and pare and then write it here. Take your time.

OpenStudy (calculusfunctions):

*paper.

OpenStudy (anonymous):

So do u find the factor of 2 tan^2 x + 5 tan x -3 ?

OpenStudy (calculusfunctions):

Not yet and where do you get 5tan x as the middle term? That's incorrect. Please try again.

OpenStudy (anonymous):

Im sorry its 2 tan^2 x + tan x -3

OpenStudy (calculusfunctions):

Yes!

OpenStudy (calculusfunctions):

Now you have\[\int\limits_{}^{}\frac{ \tan x \sec ^{2}x }{ 2\tan ^{2}x +\tan x -3 }dx\]Correct?

OpenStudy (calculusfunctions):

Agree so far?

OpenStudy (anonymous):

Yes i agree to that

OpenStudy (calculusfunctions):

Good! Now are you familiar with the substitution method?

OpenStudy (calculusfunctions):

If you are then what should you let u equal?

OpenStudy (anonymous):

Okay so now we're using the substituition method? I was thinking of the partial fraction method

OpenStudy (calculusfunctions):

Not yet! Just work with me here, alright?

OpenStudy (calculusfunctions):

So what should u equal?

OpenStudy (anonymous):

Okay. U = tan x is it?

OpenStudy (calculusfunctions):

Perfect!! Now if u = tan x then du = ?

OpenStudy (anonymous):

Sec^2 x as the du

OpenStudy (calculusfunctions):

Again Perfect!! You're on a roll. Let's not loose this momentum now. OK so now after substitution, what will the integral look like?

OpenStudy (anonymous):

Integral will look like u / (2u^2 + u -3) du

OpenStudy (calculusfunctions):

Beautiful! Now what do you propose we should do?

OpenStudy (anonymous):

The partial fraction method?

OpenStudy (calculusfunctions):

Except you forgot the little integral sign in front. lol

OpenStudy (calculusfunctions):

Not yet!

OpenStudy (anonymous):

Haha sorry. Okay i'll include that in my working later

OpenStudy (calculusfunctions):

Before we consider partial fractions, what should we do first.

OpenStudy (anonymous):

Find the factor of 2u^2 + u -3

OpenStudy (calculusfunctions):

Correct! Factor the denominator. Go ahead and tell me the factored form of that polynomial.

OpenStudy (anonymous):

(X-1) and (3x+2)

OpenStudy (calculusfunctions):

No, try again. If you expand or FOIL will you get the polynomial expression you tried to factor? I don't think so.

OpenStudy (calculusfunctions):

Please try again. I'm sure you probably made a careless error somewhere. We all do that so nothing to get discouraged over. Simply try again. Go ahead.

OpenStudy (anonymous):

Alright sorry. Its u-1 and 2u+3

OpenStudy (calculusfunctions):

Excellent! Now what?

OpenStudy (anonymous):

Now i'll break it into the a and b of the partial fraction?

OpenStudy (calculusfunctions):

Finally the partial fractions you've been wishing for all along. Great! So go ahead and tell me what you will get as your partial fractions. Take your time.

OpenStudy (anonymous):

Hahah okay thanks. I think i get it from here. So its gonna be 3/2 ln (2u+3) + 1/5 ln (x-1) then substitute u back with tan x

OpenStudy (calculusfunctions):

You mean ln (u - 1). Please be careful!!!

OpenStudy (anonymous):

Okay okay yes that

OpenStudy (calculusfunctions):

Now what do we do for the grand finale?

OpenStudy (anonymous):

Substitute the u back into tan x

OpenStudy (calculusfunctions):

You mean replace each u with tan x and also don't forget to add a constant c at the end because without it the solution is still incorrect. Alright! So please let me see your final solution.

OpenStudy (anonymous):

3/2 ln (tanx +3) + 1/5 ln (tan x -1) + c

OpenStudy (calculusfunctions):

Not quite! Almost but not quite! Please double check.

OpenStudy (anonymous):

Oh wat is wrong then?

OpenStudy (calculusfunctions):

Also I just realized that before you had (3/2) ln(2u + 3) as the integral but in fact it's (3/5)[(1/2) ln (2u + 3) which makes sense because the derivative of (1/2) ln (2u + 3) is 1/(2u + 3). Correct? Can you see where you went wrong?

OpenStudy (calculusfunctions):

You never actually told me what you got for A and B.

OpenStudy (calculusfunctions):

I should have asked you for the values of A and B before we went any further. Could you please tell me now?

OpenStudy (calculusfunctions):

Hint: Values for A and B are both rational numbers. What did you get?

OpenStudy (calculusfunctions):

Did she leave?

hartnn (hartnn):

yup. she'll return , ithink, some net problems....

OpenStudy (calculusfunctions):

OH! Thanks!

OpenStudy (calculusfunctions):

Can you please tell me when she returns? Meanwhile I can perhaps help someone else. Thanks!

hartnn (hartnn):

if i m online then i'll tell.

OpenStudy (calculusfunctions):

Thanks, actually I should get the notification also.

OpenStudy (calculusfunctions):

OK I see you're back.

OpenStudy (anonymous):

Yes sorry net problem. Got hang. I got a and b as a fraction is that correct?

OpenStudy (calculusfunctions):

So Can you tell me now what you got for A and B?

OpenStudy (anonymous):

3/2 and 1/5

OpenStudy (calculusfunctions):

I assume you mean A = 3/2 and B = 1/5. Only one of them is correct.

OpenStudy (calculusfunctions):

Please go back double check your work.

OpenStudy (anonymous):

oh really? is it 2/3 then?

OpenStudy (calculusfunctions):

Which one do you think is incorrect and no, neither of them are 2/3. This is why I said that if you show me all your steps then I can see where you may have gone wrong.

OpenStudy (calculusfunctions):

Fine I'll make it a little easy for you. B = 1/5 is correct. A is wrong.

OpenStudy (anonymous):

okay wait hold on ill figure it out

OpenStudy (calculusfunctions):

Take your time. I just want you to understand and be able to solve future problems with confidence.

OpenStudy (anonymous):

okay thanks. @calculusfunctions i need to go now. i hope u'll be there for me if i have further questions needed to be asked. i perfectly understand it now. thanks!

OpenStudy (calculusfunctions):

For you reference before you go, you should have A = 3/5 and B= 1/5. And the final answer should be\[\frac{ 3 }{ 10 }\ln (2\tan x +3)+\frac{ 1 }{ 5 }\ln (\tan x -1)+c\]Otherwise, other than that careless mistake,you did great!